# Spectral Truncation Kernels: Noncommutativity in $C^{*}$-algebraic Kernel Machines

### v1

@@ Abstract
In this paper, we propose a new class of positive definite kernels based on the spectral truncation, which has been discussed in the fields of noncommutative geometry and $C^{*}$-algebra. We focus on kernels whose inputs and outputs are functions and generalize existing kernels, such as polynomial, product, and separable kernels, by introducing a truncation parameter $n$ that describes the noncommutativity of the products appearing in the kernels. When $n$ goes to infinity, the proposed kernels tend to the existing commutative kernels. If $n$ is finite, they exhibit different behavior, and the noncommutativity induces interactions along the data function domain. We show that the truncation parameter $n$ is a governing factor leading to performance enhancement: by setting an appropriate $n$, we can balance the representation power and the complexity of the representation space. The flexibility of the proposed class of kernels allows us to go beyond previous commutative kernels.
@@ 1 Introduction
Kernel methods have been one of the most fundamental tools in machine learning [1, 2, 3, 4, 5]. They have been applied, for example, to ridge regression, principal component analysis, and support vector machine. Kernel methods are characterized by function spaces called reproducing kernel Hilbert spaces (RKHSs), which are constructed by positive definite kernels. Typical positive definite kernels include the polynomial kernel, Gaussian kernel, and Laplacian kernel. Product kernels, which are constructed by the product of multiple kernels, have also been considered [1, 6].
Standard positive definite kernels are real- or complex-valued, and are well-suited to learn scalar-valued functions. Kernel methods for vector-, function-, or operator-valued outputs have also been investigated [4, 7, 8]. The kernel, in this case, is instead an operator- or function-valued positive definite function, and the associated feature space is a vector-valued RKHS (vvRKHS) [9, 7, 10] or a reproducing kernel Hilbert $C^{*}$-module (RKHM) [8]. RKHM is a generalization of RKHS by means of $C^{*}$-algebra, and it also generalizes vvRKHS. $C^{*}$-algebra is a generalization of the space of complex values and has a structure of the norm, product, and involution [11, 12]. It unifies operators and functions. Thus, the operator- and function-valued kernels for vvRKHSs and RKHMs can be unified as $C^{*}$-algebra-valued kernels. It has been shown that the norms in $C^{*}$-algebras alleviate the dependency of the generalization bound on the output dimension [13] and the product structure of a $C^{*}$-algebra goes beyond the convolution operation [14].
@@
One big challenge for kernel methods is the choice of positive definite kernels. Especially for vvRKHSs and RKHMs, the construction of good $C^{*}$-algebra-valued kernels is not obvious. A typical kernel is the separable kernel, which is defined by the product of a real- or complex-valued kernel and a positive semi-definite operator [4]. Attempts have been made to construct positive definite kernels that go beyond separable kernels. Huusari and Kadri [15] proposed entangled kernels based on concepts from quantum computing, such as partial trace and entanglement. Hashimoto et al. [14] proposed to use the product of circulant matrices and general squared matrices to construct positive definite kernels. By using the kernel, we can generalize the convolution and capture the effect of interactions of different Fourier components.
@@
In this paper, we propose a new class of $C^{*}$-algebra-valued positive definite kernels based on the spectral truncation, which has been discussed in the field of noncommutative geometry and $C^{*}$-algebra [16, 17, 18]. We focus on the case of functional inputs and outputs, and consider the $C^{*}$-algebra of continuous functions. We generalize typical existing kernels: polynomial, product, and separable kernels using the pair of maps called the approximate order isomorphism, which takes an important role in the spectral truncation [19, 17, 20]. The proposed kernels are parameterized by a natural number $n$ that corresponds to the dimension of the truncated space. We approximate functions on the $n$-dimensional truncated space and obtain $n$ by $n$ Toeplitz matrices. Thus, $n$ describes the resolution of the discretization of the function, and we call it the truncation parameter. We show that the proposed kernels converge to the existing kernels as $n$ goes to infinity.
@@
To analyze functional data, we have been focused on how to interpolate observed data to obtain functions [21, 22, 8]. On the other hand, our results shed light on good effects of the discretization (setting $n$ as a finite number) on learning process. Indded, the truncation parameter $n$ takes an important role from at least two perspectives. First, one of the important structures in the existing kernels is the product, and $n$ describes the noncommutativity of the product. Although the product of two functions is too simple to capture structures of these functions, the product of two approximated Toeplitz matrices is more complex, which enables us to capture the structures of the functions. More precisely, for two functions $a$ and $b$, their product $ab$ is defined as just the pointwise product of each value, i.e., $ab(z)=a(z)b(z)$, and the product is commutative, i.e., $ab=ba$. Therefore, we cannot capture the relationship between $a(z_{1})$ and $b(z_{2})$ for $z_{1}\neq z_{2}$ with this product. On the other hand, the product of two Toeplitz matrices is not commutative, and it can be shown that this product involves the product of $a(z_{1})$ and $b(z_{2})$ for $z_{1}\neq z_{2}$. This enables us to capture the relationship between $a(z_{1})$ and $b(z_{2})$. Second, the parameter $n$ controls the tradeoff between the representation power and the model complexity of the RKHM. We show that if $n$ is small, then the representation power is low, and the model complexity is small. On the other hand, if $n$ is large, then the representation power is high, and the model complexity is large. The expectation loss for learning a function is decomposed into the empirical loss, controlled by the representation power, and the complexity. By introducing the parameter $n$ and setting $n$ to balance the empirical loss and the complexity, we obtain higher performance compared to the case of existing kernels.
@@
  • We propose a new class of function-valued positive definite kernels that generalize existing kernels: polynomial, product, and separable kernels. The proposed kernels are based on the spectral truncation and are indexed by a truncation parameter $n$, which allows us to introduce noncommutativity into the learning process. The proposed kernels induce the interactions along the data function domain. (Definition 6 and Subsection 3.1)
  • We derive a generalization bound for the learning problem in the RKHMs associated with the proposed kernels. The bound implies that the parameter $n$ controls the tradeoff between the representation power and the model complexity of the RKHMs. (Section 4)
  • We apply the proposed kernels to kernel ridge regression. The result illustrates the advantage of introducing the truncation parameter $n$ in practical settings. (Section 5)
We emphasize that although we focus on the case where the inputs and outputs are functions and three types of kernels, our idea of constructing new positive definite kernels is general. We establish a new framework that allows us to go beyond the usual commutative kernels and to assign a significance to noncommutative structures by discretizing functions.
@@ 2 Preliminaries
We review mathematical notions required in this paper.
@@ 2.1 $C^{*}$-algebra and Reproducing kernel Hilbert $C^{*}$-module
    $\mathcal{A}$ is a normed space endowed with $\|\cdot\|_{\mathcal{A}}$, and for $a,b\in\mathcal{A}$, $\|ab\|_{\mathcal{A}}\leq\|a\|_{\mathcal{A}}\|b\|_{\mathcal{A}}$ holds. In addition, $\mathcal{A}$ is complete with respect to $\|\cdot\|_{\mathcal{A}}$.
@@
If there exists $a\in\mathcal{A}$ such that $ab=b=ba$ for any $b\in\mathcal{A}$, $a$ is called the unit, and denoted by $1_{\mathcal{A}}$.
@@
Let $\mathbb{T}=\mathbb{R}/2\pi\mathbb{Z}$ be the torus and $C(\mathbb{T})$ be the space of continuous functions on $\mathbb{T}$. Then, $\mathcal{A}:=C(\mathbb{T})$ is a $C^{*}$-algebra by means of the product: $(cd)(z)=c(z)d(z)$, involution: $c^{*}(z)=\overline{c(z)}$, and norm: $\|c\|_{\mathcal{A}}=\sup_{z\in\mathbb{T}}|c(z)|$ for $c,d\in\mathcal{A}$.
We review basic notions regarding $C^{*}$-algebra. In the following, let $\mathcal{A}$ be a $C^{*}$-algbra.
@@
An element $a$ of $\mathcal{A}$ is called positive if there exists $b\in\mathcal{A}$ such that $a=b^{*}b$ holds. For $a,b\in\mathcal{A}$, we write $a\leq_{\mathcal{A}}b$ if $b-a$ is positive. We denote by $\mathcal{A}_{+}$ the subset of $\mathcal{A}$ composed of all positive elements in $\mathcal{A}$.
@@
$\bigg{\langle}\sum_{i=1}^{n}\phi(x_{i})c_{i},\sum_{j=1}^{l}\phi(y_{j})d_{j}\bigg{\rangle}_{k}=\sum_{i=1}^{n}\sum_{j=1}^{l}c_{i}^{*}k(x_{i},y_{j})d_{j},$
@@
The norm in $\mathcal{M}_{k,0}$ is defined as $\|f\|_{k}^{2}=\|\left\langle f,f\right\rangle_{k}\|_{\mathcal{A}}$, and we can also define the $\mathcal{A}$-valued absolute value $|\cdot|_{k}$ as $|f|_{k}^{2}=\left\langle f,f\right\rangle_{k}$. The reproducing kernel Hilbert $\mathcal{A}$-module (RKHM) $\mathcal{M}_{k}$ associated with $k$ is defined as the completion of $\mathcal{M}_{k,0}$ with respect to $\|\cdot\|_{k}$. See, for example, the references [12, 11, 23, 8] for more details about $C^{*}$-algebra and RKHM.
@@ 2.2 Spectral truncation on the torus
The product in $C(\mathbb{T})$ is commutative. However, by approximating the multiplication of a function $x\in C(\mathbb{T})$ by a Toeplitz matrix, we can obtain a noncommutative product structure. Let $e_{j}(z)=\mathrm{e}^{\mathrm{i}jz}$ for $j\in\mathbb{Z}$ and $z\in\mathbb{T}$ be the Fourier function and $M_{x}$ be the multiplication operator defined on $L^{2}(\mathbb{T})$ with respect to $x$. Here, $\mathrm{i}$ is the imaginary unit, and $L^{2}(\mathbb{T})$ is the space of square integrable functions on $\mathbb{T}$. Let $P_{n}$ be the orthogonal projection onto the $n$-dimensional subspace $\operatorname{Span}\{e_{1},\ldots,e_{n}\}$. We approximate $M_{x}$ by $P_{n}M_{x}P_{n}$, i.e., by truncating the spectrum. Then, the $(j,l)$-entry of the representation matrix $R_{n}(x)\in\mathbb{C}^{n\times n}$ of $P_{n}M_{x}P_{n}$ is written as
@@
Since the $(j,l)$-entry of $R_{n}(x)$ depends only on $j-l$, $R_{n}(x)$ is a Toeplitz matrix [24]. It is characterized only by a vector $r_{n}(x)\in\mathbb{C}^{2n-1}$, where the $(j-l)$th element of $r_{n}(x)$ is $R_{n}(x)_{j,l}$. Note that $R_{n}(x)_{j,l}$ is regarded as the $(j-l)$th Fourier component of $x$, and the vector $r_{n}(x)$ is the coordinate of the function $x$ projected on the space $\operatorname{Span}\{e_{-(n-1)},\ldots,e_{n-1}\}$.
For a matrix $A\in\mathbb{C}^{n\times n}$, let $S_{n}(A)\in C(\mathbb{T})$ be the function defined as $S_{n}(A)(z)=1/n\sum_{j,l=0}^{n-1}A_{j,l}\mathrm{e}^{\mathrm{i}(j-l)z}$, where $A_{j,l}$ is the $(j,l)$-entry of $A$. The map $S_{n}$ takes the representation matrix $R_{n}(x)$ back to a function that approximates the original function $x$. Indeed, we have
@@
where $F_{n}(t)=1/n\sum_{j=0}^{n-1}\sum_{l=-j}^{j}\mathrm{e}^{\mathrm{i}lt}$ is the Fejér kernel and $\ast$ represents the convolution. By the following proposition, for each $z\in\mathbb{T}$, $S_{n}(R_{n}(x))(z)$ converges to $x(z)$ as $n\to\infty$ [25, 17].
@@
The pair of maps $R_{n}$ and $S_{n}$ is called the approximate order isomorphism.

### v3

@@ Abstract
$C^{*}$-algebra-valued kernels could pave the way for the next generation of kernel machines. To further our fundamental understanding of learning with $C^{*}$-algebraic kernels, we propose a new class of positive definite kernels based on the spectral truncation. We focus on kernels whose inputs and outputs are vectors or functions and generalize typical kernels by introducing the noncommutativity of the products appearing in the kernels. The noncommutativity induces interactions along the data function domain. We show that it is a governing factor leading to performance enhancement: we can balance the representation power and the model complexity. We also propose a deep learning perspective to increase the representation capacity of spectral truncation kernels. The flexibility of the proposed class of kernels allows us to go beyond previous commutative kernels, addressing two of the foremost issues regarding learning in vector-valued RKHSs, namely the choice of the kernel and the computational cost.
@@ 1 Introduction
Kernel methods have been one of the most fundamental tools in machine learning (Schölkopf & Smola, 2001; Gretton et al., 2007; Hofmann et al., 2008; Muandet et al., 2017). They have been applied, for example, to ridge regression, principal component analysis, and support vector machine. Kernel methods are characterized by reproducing kernel Hilbert spaces (RKHSs), which are constructed by positive definite kernels. Typical positive definite kernels include the polynomial kernel, Gaussian kernel, and Laplacian kernel. Product kernels, which are constructed by the product of multiple kernels, have also been considered (Schölkopf & Smola, 2001; Thomas, 2008).
@@
Standard positive definite kernels are scalar-valued, and are well-suited to learn scalar-valued functions. Kernel methods for vector- and function-valued outputs have also been investigated (Álvarez et al., 2012; Kadri et al., 2016). The kernels, in these cases, are instead operator-valued, and the associated feature space is vector-valued RKHSs (vvRKHSs) (Kadri et al., 2012, 2016; Minh et al., 2016). There are at least two challenges for vvRKHS methods: the computational cost and choice of kernels. A typical kernel is the separable kernel, which is defined by the product of a scalar-valued kernel and a positive semi-definite operator (Álvarez et al., 2012). Although applying the separable kernel is computationally efficient, it identifies dependencies between input and output variables separately, and cannot reflect information of input variables properly to output variables. Several attempts have been made to construct kernels that go beyond separable kernels. Huusari & Kadri (2021) proposed entangled kernels based on concepts from quantum computing, such as partial trace and entanglement. Hashimoto et al. (2023a) proposed to use the product of circulant matrices and general squared matrices to construct kernels. Using this kernel, one can generalize the convolution and capture the effect of interactions of different Fourier components. However, if we need an $m$-dimensional vector-valued outputs with nonseparable kernels, then we have to construct an $mN$ by $mN$ Gram matrix, where $N$ is the sample size, and the computational cost is $O(m^{3}N^{3})$ in general. Thus, with vvRKHSs, we cannot resolve the two issues at the same time.
In this work, we address these two challenges by introducing a new class of kernels based on the framework of reproducing kernel Hilbert $C^{*}$-module (RKHM) (Hashimoto et al., 2021). RKHM is a generalization of RKHS by means of $C^{*}$-algebra. $C^{*}$-algebra is a generalization of the space of complex values and has structures of the norm, product, and involution (Murphy, 1990; Lance, 1995). It unifies operators and functions. In this framework, kernels are generalized to $C^{*}$-algebra-valued kernel functions and allow us to consider function-valued kernels, leading to function-valued Gram matrices. By evaluating the values of the function-valued Gram matrix at $m$ different points, we obtain $m$ scalar-valued Gram matrices. This allows us to obtain an $m$-dimensional vector-valued outputs with the computational cost of $O(mN^{3})$, which alleviates the dependency on $m$ from cubic to linear compared to the case of vvRKHSs with nonseparable kernels.
@@
To obtain kernels going beyond the separable kernels, we propose a new class of $C^{*}$-algebra-valued positive definite kernels based on the spectral truncation, which has been discussed in the fields of noncommutative geometry and $C^{*}$-algebra (D’Andrea et al., 2014; van Suijlekom, 2021; Connes & van Suijlekom, 2021). Our class of kernels is large and encompasses previous commutative kernel functions such as polynomial, product, and separable kernels. The proposed kernels are parameterized by a natural number $n$ corresponding to the dimension of the truncated space. They can be applied to both vector and functional inputs. For vector inputs, we regard the elements of them as the values of functions. We approximate the input functions on the $n$-dimensional truncated space and obtain $n$ by $n$ Toeplitz matrices, whose $(i,j)$-entry depends only on $i-j$. Thus, $n$ describes the resolution of the discretization, and we call it the truncation parameter. We show that the proposed kernels converge to the commutative kernels as $n$ goes to infinity.
@@
In the setting of functional data (Jim Ramsay, 2005; Wang et al., 2016; Hashimoto et al., 2021), the proposed truncation kernels shed light on the good effects of discretization (setting $n$ as a finite number) on the learning process. The truncation parameter $n$ plays an important role from at least two perspectives. First, an important structure in typical commutative kernels is the product, and $n$ describes the noncommutativity of the product. For two functions $a$ and $b$, their product $ab$ is defined as just the pointwise product of each value, i.e., $ab(z)=a(z)b(z)$, and the product is commutative, i.e., $ab=ba$. Therefore, we cannot capture the relationship between $a(z_{1})$ and $b(z_{2})$ for $z_{1}\neq z_{2}$ with this product. This situation corresponds to the case of $n=\infty$. On the other hand, the product of two Toeplitz matrices is not commutative, and it can be shown that this product involves the product of $a(z_{1})$ and $b(z_{2})$ for $z_{1}\neq z_{2}$. This enables us to capture the relationship between $a(z_{1})$ and $b(z_{2})$. This corresponds to the case $n<\infty$. Second, the parameter $n$ controls the tradeoff between the representation power and the model complexity. We show that if $n$ is small, then the representation power is low, and the model complexity is small. On the other hand, if $n$ is large, then the representation power is high, and the model complexity is large. The expectation loss for learning a function is controlled by the representation power and the complexity. By introducing the parameter $n$ and setting $n$ to balance them, we obtain higher performance compared to the typical commutative kernels. In addition, to resolve the issue of the interactions becoming weaker as $n$ becomes larger (i.e., representation power becomes higher), we propose a deep approach to obtain models with higher representation powers while keeping the interactions.
@@
  • We propose spectral truncation kernels, a new class of function-valued positive definite kernels that generalize typical kernels. The proposed kernels are based on the spectral truncation and are indexed by a truncation parameter, which induces the interactions along the data function domain by introducing the noncommutativity into the learning process. (see Definition 6 and Subsection 3.1)
  • We derive a generalization bound for the learning problem in the RKHMs associated with the proposed kernels. The bound implies that the parameter $n$ controls the tradeoff between the representation power and the model complexity. (see Section 4)
  • We propose a deep approach to gain the representation power while keeping the noncommutativity in the kernels. (see Section 6)
Our framework allows us to go beyond the usual commutative kernels and to develop a new perspective on the significance of noncommutative structures in learning systems.
@@ 2.1 $C^{*}$-algebra and Reproducing kernel Hilbert $C^{*}$-module
    $\mathcal{A}$ is a normed space endowed with a norm $\|\cdot\|_{\mathcal{A}}$, and for $a,b\in\mathcal{A}$, $\|ab\|_{\mathcal{A}}\leq\|a\|_{\mathcal{A}}\|b\|_{\mathcal{A}}$ holds. In addition, $\mathcal{A}$ is complete with respect to $\|\cdot\|_{\mathcal{A}}$.
@@
If there exists $a\in\mathcal{A}$ such that $ab=b=ba$ for any $b\in\mathcal{A}$, $a$ is called the unit and denoted by $1_{\mathcal{A}}$.
@@
Let $\mathbb{T}=\mathbb{R}/2\pi\mathbb{Z}$ be the torus and $C(\mathbb{T})$ be the space of continuous functions on $\mathbb{T}$. Then, $\mathcal{A}:=C(\mathbb{T})$ is a $C^{*}$-algebra by means of the product: $(cd)(z)=c(z)d(z)$, involution: $c^{*}(z)=\overline{c(z)}$, and norm: $\|c\|_{\mathcal{A}}=\sup_{z\in\mathbb{T}}|c(z)|$ for $c,d\in\mathcal{A}$. The unit is the constant function $1_{\mathcal{A}}\equiv 1$.
We now review basic notions regarding $C^{*}$-algebra. In the following, let $\mathcal{A}$ be a $C^{*}$-algbra.
@@
An element $a$ of $\mathcal{A}$ is called positive if there exists $b\in\mathcal{A}$ such that $a=b^{*}b$ holds. For $a,b\in\mathcal{A}$, we write $a\leq_{\mathcal{A}}b$ if $b-a$ is positive.
@@
$\bigg{\langle}\sum_{i=1}^{N}\phi(x_{i})c_{i},\sum_{j=1}^{M}\phi(y_{j})d_{j}\bigg{\rangle}_{k}=\sum_{i=1}^{N}\sum_{j=1}^{M}c_{i}^{*}k(x_{i},y_{j})d_{j},$
@@
The norm in $\mathcal{M}_{k,0}$ is defined as $\|f\|_{k}^{2}=\|\left\langle f,f\right\rangle_{k}\|_{\mathcal{A}}$, and we can also define the $\mathcal{A}$-valued absolute value $|\cdot|_{k}$ as $|f|_{k}^{2}=\left\langle f,f\right\rangle_{k}$. The reproducing kernel Hilbert $\mathcal{A}$-module (RKHM) $\mathcal{M}_{k}$ associated with $k$ is defined as the completion of $\mathcal{M}_{k,0}$ with respect to $\|\cdot\|_{k}$. See, for example, the references (Lance, 1995; Murphy, 1990; Hashimoto et al., 2021) for more details about $C^{*}$-algebra and RKHM.
@@ 2.2 Spectral truncation on the torus
The product in $C(\mathbb{T})$ is commutative. However, by approximating the multiplication of a function $x\in C(\mathbb{T})$ by a Toeplitz matrix, we can obtain a noncommutative product structure. Let $e_{j}(z)=\mathrm{e}^{\mathrm{i}jz}$ for $j\in\mathbb{Z}$ and $z\in\mathbb{T}$ be the Fourier function and $M_{x}$ be the multiplication operator defined on $L^{2}(\mathbb{T})$ with respect to $x$. Here, $\mathrm{i}$ is the imaginary unit, and $L^{2}(\mathbb{T})$ is the space of square-integrable functions on $\mathbb{T}$. Let $P_{n}$ be the orthogonal projection onto the $n$-dimensional subspace $\operatorname{Span}\{e_{1},\ldots,e_{n}\}$. We approximate $M_{x}$ by $P_{n}M_{x}P_{n}$, i.e., by truncating the spectrum. Then, the $(j,l)$-entry of the representation matrix $R_{n}(x)\in\mathbb{C}^{n\times n}$ of $P_{n}M_{x}P_{n}$ is written as
@@
Since the $(j,l)$-entry of $R_{n}(x)$ depends only on $j-l$, $R_{n}(x)$ is a Toeplitz matrix (Gray, 2006). It is characterized only by a vector $r_{n}(x)\in\mathbb{C}^{2n-1}$, where the $(j-l)$th element of $r_{n}(x)$ is $R_{n}(x)_{j,l}$. Note that $R_{n}(x)_{j,l}$ is regarded as the $(j-l)$th Fourier component of $x$, and the vector $r_{n}(x)$ is the coordinate of the function $x$ projected on the space $\operatorname{Span}\{e_{-(n-1)},\ldots,e_{n-1}\}$.
For a matrix $A\in\mathbb{C}^{n\times n}$, let $S_{n}(A)\in C(\mathbb{T})$ be the function defined as $S_{n}(A)(z)=(1/n)\sum_{j,l=0}^{n-1}A_{j,l}\mathrm{e}^{\mathrm{i}(j-l)z}$, where $A_{j,l}$ is the $(j,l)$-entry of $A$. The map $S_{n}$ takes the representation matrix $R_{n}(x)$ back to a function that approximates the original function $x$. Indeed, we have
@@
where $F_{n}(t)=(1/n)\sum_{j=0}^{n-1}\sum_{l=-j}^{j}\mathrm{e}^{\mathrm{i}lt}$ is the Fejér kernel and $\ast$ represents the convolution. By the following proposition, for each $z\in\mathbb{T}$, $S_{n}(R_{n}(x))(z)$ converges to $x(z)$ as $n\to\infty$ (Brandolini & Travaglini, 1997; van Suijlekom, 2021).
@@
A generalization of the Fejér kernel on $\mathbb{T}$ to that on $\mathbb{T}^{q}$ with respect to the sum over a polyhedron has been theoretically investigated (Travaglini, 1994; Brandolini & Travaglini, 1997). Let $P$ be a polyhedron and $jP=\{jz\,\mid\,z\in P\}$ for $j\in\mathbb{Z}$. For $t\in\mathbb{T}^{q}$, the Fejér kernel on $\mathbb{T}^{q}$ is defined as
$F_{n}^{q,P}(t)=\frac{1}{n}\sum_{j=1}^{n-1}\sum_{r\in jP\bigcap\mathbb{Z}^{q}}\mathrm{e}^{\mathrm{i}r\cdot t}.$ (1)